# Flat band driven itinerant magnetism in the Co-pnictides (La,Ca)Co2(As,P)2

### v1

@@ Abstract
Flat bands can induce strong electron correlation effects that help stabilize both magnetic and superconducting states. Here, we carry out angle-resolved photoemission spectroscopy and density functional theory calculations to study the electronic structure of the Co-pnictides CaCo2As2 and LaCo2P2. We find that, while the kz Fermi topology of ferromagnetic LaCo2P2 is markedly 2-dimensional, antiferromagnetic CaCo2As2 develops a 3D Fermi surface described by a zig-zag-like band dispersion perpendicular to the Co-As plane. Furthermore, the magnetism is driven by the electronic correlations of the flat bands with dxy and d orbital character at the Fermi level. Our results link the electronic dimensionality and the magnetic order, and emphasize the critical role of the As-As and P-P bond strength along the c-direction to understand the electronic band structure and the rich phase diagram of transition metal pnictides.
The unique interplay between electron-electron correlations and the high density of states at the Fermi level results in strong interactions that throw off their balance and favor unconventional orders. In the case of weakly dispersive flat bands, Coulomb repulsion dominates over the kinetic energy, giving rise to enhanced correlations especially in 2-dimensional (2D) materials and kagome metals that favor magnetism, fractional quantum Hall states and competing phases such as superconductivity and charge density waves [1, 2, 3, 4, 5, 6, 7, 8].
@@
On the other hand, despite the magnetism reported experimentally in Co-pnictides, the link between magnetic correlations and electronic dimensionality has not been hitherto addressed. The microscopic origin of the magnetic orders is particularly appealing, since Density Functional Theory (DFT) calculations unveil the presence of a flat band along the -M direction and high density of states (DOS) at the Fermi level. This debate is further stimulated by the observation of flat band induced magnetism in the sibling transition metal dichalcogenide ACo2Se2 compound [30, 31] and the coexistence of both AFM and FM spin fluctuations in SrCo2As2 [32].
@@
Here, we use angle-resolved photoemission spectroscopy (ARPES) and DFT calculations to accurately describe the orbital contribution of the band structure of AFM and FM Co-pnictides CaCo2As2 and LaCo2P2. We find that the electronic structure of CaCo2As2 is 3D, in contrast to the 2D observed in the itinerant ferromagnet LaCo2P2. Besides, we demonstrate that the magnetic moment is dominated by the dxy and d orbitals of Co that form a flat band along the - direction. Our work bring an important piece of information to understand the intertwining of electronic dimensionality and magnetism of a large family of more than 500 layered AT2X2-type compounds.
Single crystals of CaCo2As2 and LaCo2P2 were grown by Sn-flux method, as previously reported elsewhere [33, 15, 16, 34]. x-ray absorption spectroscopy (XAS) and x-ray magnetic circular dichroism (XMCD) experiments at the Co -edge were carried out at the BOREAS beamline at the ALBA synchrotron. ARPES experiments were performed at the BLOCH (MAX IV), ESM (NSLS) and LOREA (ALBA) beamlines, respectively, with experimental energy and angular resolution of and , respectively. The samples were in-situ cleaved and the measurements were taken at 20 K and above the transition temperatures.
@@
The Co-pnictides A (A = ) exhibit a body-centered tetragonal ThCr2Si2 structure (, 139 space group) characterized by alternating stacks of tetrahedral layers (Fig. 1(a)), with the strong bonding within the tetrahedra dominating the interactions, akin to the chemical structure of Fe-pnictides [35]. The conventional unit cell lattice parameters are a=b=4.01 Å and = 10.38 Å for CaCo2As2 and a=b= 3.80 Å and = 11.01 Å for LaCo2P2, with the ratio /a=2.59 and 2.9, respectively, suggesting a structural link with the magnetic order [16].
@@
Figs. 1 (c-d) display the x-ray absorption (XAS) spectra corresponding to the Co L3 ( eV) transition that already hints at a different electronic structure. The XAS lineshape strongly depends on the multiplet structure, mainly given by the Co 3d3d and 2p3d Coulomb and exchange interactions, the local crystal fields and the hybridization with the ligands [36]. The absorption edges are separated by 16 eV from the L2 edge ( eV) that result from the spin-orbit split 2p3d transitions. The Co L3 of CaCo2As2 presents five fine structures: a ‘pre-edge’ transition at 776 eV (P1), a major absorption edge with 2 main peaks at 777.8 eV (P2) and 778.2 eV (P3), two high energy shoulders at 779 eV (P4) and 782 eV (P5), characteristic of a Co2+ oxidation state, see Fig. 1 (c), and a broad absorption line corresponding to the transitions into the continuum above 786 eV. On the other hand, the L3 structure of LaCo2P2 is composed of one main peak (P1 at 779.4 eV), two shoulders at 781 eV (P2) and 784.5 eV (P3), Fig. 1 (d), suggesting a Co oxidation state higher than 2+ (3+ according to the electron counting approach) and the continuum of transitions 10 eV above the main edge [37, 38].
@@
The calculated DOS shows a continuum of metallic states, Fig. 2. In the paramagnetic (PM) calculation, we observe signatures of hybridization between Co and As 4p states, which contributes nearly 8% to the DOS of CaCo2As2, and the Ca participates with almost 3% of states in the vicinity of the Fermi level, hence, the total DOS is mainly derived from the Co 3d orbitals that develop a sharp peak at EF, Fig. 2(a). Interestingly, the sharp peak of the DOS in the PM calculation is split 0.5 eV around the Fermi level (Fig. 2(b)), suggesting that electron correlations are responsible for the magnetic order. Similarly, the DOS of LaCo2P2, Fig. 2(f), presents a sharp peak near the Fermi level ( states/eV) in the PM phase, that splits into spin polarized minority and majority bands in the FM phase due to the Stoner mechanism, Fig. 2(g).
@@
In Figs. 2(c) and (h), we show the kx-kz Fermi surface of CaCo2As2 and LaCo2P2, obtained by varying the incident photon energy from 50 to 120 eV (E=2 eV). The high symmetry points of CaCo2As2, labeled as and , are located at 78 and 105 eV, respectively, Fig. 2(d-e). The zig-zag dispersion of the bands, superimposed with the Fermi momenta (kf) obtained from the momentum distribution curves (MDC) of the bands crossing the Fermi level along the c-direction, demonstrates a 3D character of its Fermi surface, 2(c). On the other hand, the kz dependence of LaCo2P2 displays two dispersionless bands, identified as 1 and 2 in Fig. 2(h), revealing a pronounced 2-dimensionality of its electronic structure, consistent with the high magnetic field transport data reported in the literature [39].
@@
Figures 2(d-e) and 2(i-j) display the low temperature kx-ky Fermi surface (FS) of CaCo2As2 and LaCo2P2 at kz=0 and planes, being the lattice parameter of the primitive unit cell (distance between two Co planes), corresponding to the and planes. The constant energy contours of the electronic structure present C4 rotational symmetry, with a rectangular electron pocket at the BZ center [40, 41] and 4 square electron pockets at the corners, and A points, as expected from the crystal space group.
@@
Centering our attention on the kz=0 plane of CaCo2As2, Fig. 3 (a-c) displays the electronic energy-momentum band dispersion along the high-symmetry path X--M and its comparison with the projected orbital character of Co obtained ab initio in the paramagnetic state (150 K). At , there is an electron pocket derived from the dxy orbital that develops more spectral weight with LV polarized light. Along the path, we observe another electron pocket () with dxz character, which is only visible with LH polarized light. In addition to these two electron pockets, our high-temperature valence band (VB) spectra resolve a flat band (FB in Fig. 3(a) and suppl. inf. Fig. S7) along the -M direction, resulting from the hybridization of dxy and d orbitals. The momentum spread of the flat band is responsible for the high DOS at EF in Fig. 2(a) and contributes with 0.39 (dxy) and 0.12 (d) to the total magnetization, as calculated ab initio. In the plane, there is an electron pocket centered at with orbital character, visible with LH light, Fig. 3(e). Along the high-symmetry line, we can further resolve 2 highly-dispersive bands with d and dxz orbital characters that cross EF at Å-1 and 0.65Å-1, highlighted with vertical arrows in Fig. 3(e). In addition, the small electron pocket at A is also well reproduced by our DFT calculations. Nevertheless, despite the large agreement between ARPES and DFT, we note a mismatch of 0.2 eV of the broad ‘’ band at 0.4 eV below the EF, mostly identified with the dxz, dyz and d orbitals at both kz=0 and planes. We attribute this discrepancy to the magnetic frustration of CaCo2As2 [22] that is not fully captured by our DFT calculations. Indeed, the position of the ‘’ band is well reproduced by the FM calculation, see suppl. inf. Fig. S4, further supporting the frustrated square lattice scenario. At low temperature, Fig. 3(g-h), the band structure of CaCo2As2 is substantially modified. Notably, the electron pocket at disappears, an electron pocket is visible along the -X path, Fig. 3(h) (dxy and dxz hybridized bands). Besides, along the direction, we identify three highly-dispersive bands with Fermi momenta, Å-1, 0.6 Å-1 and 0.8 Å-1, marked by arrows in Fig. 3(h) (see suppl. inf. Fig. S5, for higher resolution VBs), also well reproduced by the AFM DFT calculations in Fig. 3(i). In the plane, the bands along the path have Fermi momenta Å-1, 0.6 Å-1 and 0.8 Å-1, similar to the highly-dispersive bands in the plane. These highly-disperse bands and the ‘’-shaped set of bands 0.4 eV below EF retain their orbital characters with respect to the high temperature paramagnetic phase, Figs. 3(g-h) and (j-k).
@@
The experimental band structure of the PM (150 K) and FM (20 K) phases of LaCo2P2 and the ab-initio calculations are plotted in Fig. 4. At high temperature, the ‘’ shaped band shows up at around 0.4 eV below EF, Fig. 4(a), and its orbital contribution, dxz, and d, is well discerned in the polarization dependent VB spectra. The top of this band approaches the electron pocket () and crosses EF at Å-1, vertical arrow in Fig. 4(a), with most of its spectral weight derived from the dxy orbital. In addition, the VBs with LH light show an electron-like band (), and orbital character, centered at with its bottom at a binding energy of . In the plane, Fig. 4(d-e), the electronic energy bands show similar features as in the plane, with the faint electron pocket () and electron-like band () centered at . Moreover, two highly linear dispersive bands (), visible for LV polarized light, are shown along the direction, in agreement with the DFT calculations. At low temperature, Fig. 4(g-l), we resolve a large electron pocket around and kz0, Fig. 4(g), which presents mostly d and dyz character. Such electron pocket cuts EF at Å-1 along and paths, with the dispersive band () along direction derived from the d orbital, Fig. 4(g). The VBs show linear dispersing bands along the R-Z-A path with different orbital characters, whose cuts at the EF are highlighted by arrows in Fig. 4(j) and (k), again well captured by our DFT calculations. Finally, the flat-ish band along the Z-A path, derived also from the dxy and d orbitals of Co, contributes with 0.62 and 0.22 , respectively, to the total magnetic moment, see suppl. inf. Table I. The detailed electronic structure shows a small contribution of the pnictide orbitals, with the 3p electrons of P and the 4p electrons of As contributing to the p-d bonding at higher binding energies.
I Discussion
@@
As sibling compounds of superconducting iron pnictides, CaCo2As2 and LaCo2P2 present very different band structure. 122 (Ca,Ba)Fe2As2 exhibits a second-order-like stripe pattern followed by a first-order structural transition to a cT phase at low temperature that underscores the emergence of a broken rotational symmetry state [9]. From the electronic structure point of view, the iron pnictides have several hole pockets at that nest the electron pockets around M point, driving the AFM order with a propagation vector (-0). As a direct consequence of adding one more electron from Co, the upward shift of the chemical potential by 1 eV [42] and the presence of electron pockets away from the X position in CaCo2As2 prevent the in-plane FS nesting. Interestingly, the ARPES data we report here show parallels to the dimensionality crossover observed in the 122 family. In particular, a 3D kz dispersion of the electronic band structure develops in the frustrated AFM CaCo2As2, akin to the three dimensional topology crossover observed in the orthorhombic phase at low temperature in 122 CaFe2As2 [43], yet the spin polarization changes from out to in-plane. Nevertheless, the cT phase of Fe pnictides is non-magnetic as a consequence of a major topological change of the Fermi surface that reduces the nesting between the and M pockets. Finally, the FM order in LaCo2P2 reconstructs the electronic structure and the number of dispersive bands at low temperature. This is attributed to spin splitting caused by the Stoner-driven magnetic ordering, similar to the electronic structure of RbCo2Se2 that shares the same lattice structure [44].
@@
In summary, we have presented a comprehensive analysis of the band structure that demonstrate a correlation between magnetic order, chemical pressure and electronic dimensionality in the 122 Co-pnictides CaCo2As2 and LaCo2P2. We have reported a pronounced 2-dimensionality of the electronic structure of LaCo2P2, in contrast to the 3D character of CaCo2As2. The experimental and calculated band structure indicate that Co-pnictides undergo a large Fermi-surface reconstruction at the magnetic transitions, which are are a consequence of the flat-ish bands of Co. Moreover, we have shown that the shorter As-As compared with P-P distance enhances hopping between pnictogen and transition metal and, hence favoris the the AFM ordering in CaCo2As2. Our findings deepen in the fundamental understanding of the electronic band structure of a large family of compounds that host a variety of collective ground states and provide evidence that the structural, electronic and magnetic properties are strongly entangled.
II Acknowledgements

### v2

# Flat band driven itinerant magnetism in the Co-pnictides CaCo2As2 and LaCo2P2

@@ Abstract
Flat bands can induce strong electron correlation effects that help stabilize both magnetic and superconducting states. Here, we carry out angle-resolved photoemission spectroscopy and density functional theory calculations to study the electronic structure of the Co-pnictides CaCo2As2 and LaCo2P2. We find that the kz Fermi topology of ferromagnetic LaCo2P2 is markedly 2-dimensional, while the antiferromagnetic CaCo2As2 develops a 3D Fermi surface described by a zig-zag-like band dispersion perpendicular to the Co-As plane. Furthermore, the magnetism is driven by the electronic correlations of the flat bands with dxy and d orbital character at the Fermi level. Our results link the electronic dimensionality and the magnetic order, and emphasize the critical role of the As-As and P-P bond strength along the c-direction to understand the electronic band structure and the rich phase diagram of transition metal pnictides.
The unique interplay between electron-electron correlations and the high density of states at the Fermi level results in strong interactions that throw off their balance and favor unconventional orders. In the case of weakly dispersive flat bands, Coulomb repulsion dominates over the kinetic energy, giving rise to enhanced correlations especially in 2-dimensional (2D) materials and kagome metals that favor magnetism, fractional quantum Hall states and competing phases such as superconductivity, enhanced d0 magnetism and charge density waves [1, 2, 3, 4, 5, 6, 7, 8, 9].
@@
On the other hand, despite the magnetism reported experimentally in Co-pnictides, the link between magnetic correlations, chemical structure, and electronic dimensionality/band structure has not been hitherto addressed. The microscopic origin of the magnetic orders is particularly appealing, since Density Functional Theory (DFT) calculations unveil the presence of a flat band along the -M direction and high density of states (DOS) at the Fermi level, which potentially would give rise to magnetic order through the Stoner mechanism. This debate is further stimulated by the observation of flat band induced magnetism in the sibling transition metal dichalcogenide ACo2Se2 compound [31, 32] and the coexistence of both AFM and FM spin fluctuations in SrCo2As2 [33].
@@
Here, we use angle-resolved photoemission spectroscopy (ARPES) and DFT calculations to link the chemical structure, Fermi surface topology, flat bands, and magnetic order in ACo2X2 (A= La, Ca, and X = P, AS). We find that the electronic structure of LaCo2P2 is 2D, in contrast to the 3D observed in the itinerant antiferromagnet CaCo2As2. The 3D Fermi surface topology in CaCo2As2 favors the interlayer antiferromagnetic order, provided that the As-As distance, dAs-As, is shorter than dP-P in LaCo2P2. From the DFT calculations, we achieve an accurate description of the orbital contribution to the band structure and demonstrate that the magnetic moment is dominated by the dxy and d orbitals of Co, that form a flat band along the - direction. Our work brings an important piece of information to understand the intertwining of electronic dimensionality and the magnetic order in the family of layered AT2X2-type compounds with flat band driven magnetism.
Single crystals of CaCo2As2 and LaCo2P2 were grown by Sn-flux method, as previously reported elsewhere [34, 16, 17, 35]. x-ray absorption spectroscopy (XAS) and x-ray magnetic circular dichroism (XMCD) experiments at the Co -edge were carried out at the BOREAS beamline at the ALBA synchrotron. ARPES experiments were performed at the BLOCH (MAX IV), ESM (NSLS-II), and LOREA (ALBA) beamlines, respectively, with experimental energy and angular resolution of and . The samples were in-situ cleaved and the measurements were taken at 20 K and above the transition temperatures. DFT calculations of the electronic band structure were carried out using the Vienna Ab-initio Simulation Package (VASP) software[36, 37] for the experimental structures available on Materials Project [38]. We chose the Perdew-Burke-Ernzenhof (PBE) implementation of the Generalized Gradient Approximation (GGA) for the exchange-correlation functional [39] together with Projector Augmented Wave (PAW) pseudo-potentials [40]. The -point grid was adjusted to a Gamma-centered, regular Monkhorst-Pack point set of dimensions , and the plane wave basis energy-cutoff was fixed to 600 eV. To accurately describe the d valence electrons, a series of DFT+U calculations in the rotational-invariant implementation [41] were performed to fit the optimal value of the repulsive U parameter in the d subspace to reproduce the reported experimental magnetic moment [22, 42]. It was determined that U=1 for AFM and U=1 for FM , with resulting magnetic moments of and per atom, respectively. A Wannier function basis accurately describing the band structure within eV from the Fermi level was constructed using Wannier90 [43]. Furthermore, we used the interpolated tight-binding Hamiltonian obtained in this basis to compute the Fermi surface and surface states using WannierTools [44, 45].
@@
The Co-pnictides CaCo2As2 and LaCo2P2 exhibit a body-centered tetragonal ThCr2Si2 structure (space group , no. 139) characterized by alternating stacks of tetrahedral layers (Fig. 1(a)), with the strong bonding within the tetrahedra dominating the interactions, akin to the chemical structure of Fe-pnictides [46]. The conventional unit cell lattice parameters are a = b = 4.01 Å and = 10.38 Å for CaCo2As2 and a= b = 3.80 Å and = 11.01 Å for LaCo2P2, with the ratio /a = 2.59 and 2.9, respectively, suggesting a structural link with the magnetic order [17].
@@
Figs. 1 (c-d) displays the X-ray absorption (XAS) spectra corresponding to the Co L3 ( eV) transition that already hints at a different electronic structure. The XAS lineshape strongly depends on the multiplet structure, mainly given by the Co 3d3d and 2p3d Coulomb and exchange interactions, the local crystal fields, and the hybridization with the ligands [47]. The absorption edges are separated by 16 eV from the L2 edge ( eV) that results from the spin-orbit split 2p3d transitions. The Co L3 spectrum of CaCo2As2 presents five fine structures: a ‘pre-edge’ transition at 776 eV (P1), a major absorption edge with 2 main peaks at 777.8 eV (P2) and 778.2 eV (P3), two high energy shoulders at 779 eV (P4) and 782 eV (P5), characteristic of a Co2+ oxidation state, see Fig. 1 (c), and a broad absorption line corresponding to the transitions into the continuum above 786 eV. On the other hand, the L3 spectrum of LaCo2P2 is composed of a main peak (P1 at 778.4 eV) and two shoulders at 779 eV (P2) and 782.5 eV (P3), Fig. 1 (d), suggesting the continuum of transitions 10 eV above the main edge [48, 49].
@@
The calculated DOS of CaCo2As2 shows a continuum of metallic states, see Fig. 2. In the paramagnetic (PM) calculation, we observe signatures of hybridization between Co and As 4p states, which contributes nearly 8% to the DOS of CaCo2As2, and the Ca participates with almost 3% of states in the vicinity of the Fermi level, hence, the total DOS is mainly derived from the Co 3d orbitals that develop a sharp peak at EF, Fig. 2(a). Interestingly, the sharp peak of the DOS in the PM calculation is split 0.5 eV around the Fermi level (Fig. 2(b)), suggesting that electron correlations are responsible for the magnetic order. Similarly, the DOS of LaCo2P2, Fig. 2(f), presents a sharp peak near the Fermi level ( states/eV) in the PM phase, that splits into spin-polarized minority and majority bands in the FM phase due to the Stoner mechanism, Fig. 2(g) [50].
@@
In Figs. 2(c) and (h), we show the kx-kz Fermi surface of CaCo2As2 and LaCo2P2, obtained by varying the incident photon energy from 50 to 120 eV (E=2 eV). The high symmetry points of CaCo2As2, labeled as and , are located at 78 and 105 eV, respectively, Fig. 2(d-e). The zig-zag dispersion of the bands, superimposed with the Fermi momenta (kf) obtained from the momentum distribution curves (MDC) of the bands crossing the Fermi level along the c-direction, demonstrates a 3D character of its Fermi surface of CaCo2As2, Fig. 2(c). On the other hand, the kz dependence of LaCo2P2 displays two dispersionless bands, identified as 1 and 2 in Fig. 2(h), revealing a pronounced 2-dimensionality of its electronic structure, consistent with the high magnetic field transport data reported in the literature [51].
@@
Figures 2(d-e) and 2(i-j) display the low temperature kx-ky Fermi surface of CaCo2As2 and LaCo2P2 at kz=0 and planes, being the lattice parameter of the primitive unit cell (distance between two Co planes), corresponding to the and planes. The constant energy contours of the electronic structure present C4 rotational symmetry, with a rectangular electron pocket at the BZ center [52, 53] and 4 square electron pockets at the corners, and A points, as expected from the crystal space group.
@@
Centering our attention on the kz=0 plane of CaCo2As2, Fig. 3 (a-c) displays the electronic energy-momentum band dispersion along the high-symmetry path X--M and its comparison with the projected orbital character of Co obtained ab initio in the paramagnetic state (150 K). At , there is an electron pocket derived from the dxy orbital that develops more spectral weight with LV polarized light. Along the path, we observe another electron pocket () with dxz character, which is only visible with LH polarized light. In addition to these two electron pockets, our high-temperature valence band (VB) spectra resolve a flat band (FB in Fig. 3(a) and suppl. inf. Fig. S7) along the -M direction, resulting from the hybridization of dxy and d orbitals. The momentum spread of the flat band is responsible for the high DOS at EF in Fig. 2(a) and contributes 0.39 (dxy) and 0.12 (d) to the total magnetization, as calculated ab initio. In the plane, there is an electron pocket centered at with orbital character, visible with LH light, Fig. 3(e). Along the high-symmetry line, we can further resolve 2 highly-dispersive bands with d and dxz orbital characters that cross EF at Å-1 and 0.65Å-1, highlighted with vertical arrows in Fig. 3(e). In addition, the small electron pocket at A is also well reproduced by our DFT calculations. Nevertheless, despite the good agreement between ARPES and DFT, we note a mismatch of 0.2 eV of the broad ‘’ band at 0.4 eV below the EF, mostly identified with the dxz, dyz and d orbitals in both kz=0 and planes. We attribute this discrepancy to the magnetic frustration of CaCo2As2 [23] that is not fully captured by our DFT calculations. Indeed, the position of the ‘’ band is well reproduced by the FM calculation, see suppl. inf. Fig. S4 further supports the frustrated square lattice scenario. At low temperature, Fig. 3(g-h), the band structure of CaCo2As2 is substantially modified. Notably, the electron pocket at disappears, an electron pocket is visible along the -X path, Fig. 3(h) (dxy and dxz hybridized bands). Besides, along the direction, we identify three highly-dispersive bands with Fermi momenta, Å-1, 0.6 Å-1 and 0.8 Å-1, marked by arrows in Fig. 3(h) (suppl. inf. Fig. S7, for higher resolution VBs), also well reproduced by the AFM DFT calculations in Fig. 3(i). In the plane, the bands along the path have Fermi momenta Å-1, 0.6 Å-1 and 0.8 Å-1, similar to the highly-dispersive bands in the plane. These highly-disperse bands and the ‘’-shaped set of bands 0.4 eV below EF retain their orbital characters with respect to the high temperature paramagnetic phase, Figs. 3(g-h) and (j-k).
@@
The experimental band structure of the PM (150 K) and FM (20 K) phases of LaCo2P2 and the ab-initio calculations are plotted in Fig. 4. At high temperature, the ‘’ shaped band shows up at around 0.4 eV below EF, Fig. 4(a), and its orbital contributions, dxz, and d, are well discerned in the polarization dependent VB spectra. The top of this band approaches the electron pocket () and crosses EF at Å-1, vertical arrow in Fig. 4(a), with most of its spectral weight derived from the dxy orbital. In addition, the VBs with LH light show an electron-like band (), and orbital character, centered at with its bottom at a binding energy of . In the plane, Fig. 4(d-e), the electronic energy bands show similar features as in the plane, with the faint electron pocket () and electron-like band () centered at . Moreover, two highly linear dispersive bands (), visible for LV polarized light, are seen along the direction, in agreement with the DFT calculations. At low temperature, Fig. 4(g-l), we resolve a large electron pocket around and kz0, Fig. 4(g), which presents mostly d and dyz character. Such electron pocket cuts EF at Å-1 along and paths, with the dispersive band () along direction derived from the d orbital, Fig. 4(g). The VBs show linear dispersing bands along the R-Z-A path with different orbital characters, whose cuts at the EF are highlighted by arrows in Fig. 4(j) and (k), again well captured by our DFT calculations. Finally, the flat-ish band along the Z-A path is derived also from the dxy and d orbitals of Co, which contribute 0.62 and 0.22 , respectively, to the total magnetization, see suppl. inf. Table I. The detailed electronic structure show small contributions of the pnictide orbitals, with the 3p electrons of P and the 4p electrons of As contributing to the p-d bonding at higher binding energies.
@@
As sibling compounds of superconducting iron pnictides, CaCo2As2 and LaCo2P2 present very different band structures. (Ca,Ba)Fe2As2 exhibits a second-order-like stripe pattern followed by a first-order structural transition to a cT phase at low temperature that underscores the emergence of a broken rotational symmetry state [10]. From the electronic structure point of view, the iron pnictides have several hole pockets at that nest the electron pockets around the M point, driving the AFM order with a propagation vector (-0). As a direct consequence of adding one more electron from Co, the upward shift of the chemical potential by 1 eV [54] and the presence of electron pockets away from the X position in CaCo2As2 prevent the in-plane Fermi surface nesting. Interestingly, the ARPES data we report here show parallels to the dimensionality crossover observed in the AFe2As2 family. In particular, a 3D kz dispersion of the electronic band structure develops in the frustrated AFM CaCo2As2, akin to the three-dimensional topology crossover observed in the orthorhombic phase at low temperature in CaFe2As2 [55], yet the spin polarization changes from out-of-plane to in-plane. Nevertheless, the cT phase of Fe pnictides is non-magnetic as a consequence of a major topological change of the Fermi surface that reduces the nesting between the and M pockets. Finally, the FM order in LaCo2P2 reconstructs the electronic structure and the number of dispersive bands at low temperature. This is attributed to spin splitting caused by the Stoner-driven magnetic ordering [50], similar to the electronic structure of RbCo2Se2 that shares the same lattice structure [32].
@@
In summary, we have presented a comprehensive analysis of the band structure that demonstrates a correlation between magnetic order, chemical pressure, and electronic dimensionality in the Co-pnictides CaCo2As2 and LaCo2P2. We have reported a pronounced 2-dimensionality of the electronic structure of LaCo2P2, in contrast to the 3D character of CaCo2As2. The experimental and calculated band structure indicate that these pnictides undergo a large Fermi-surface reconstruction at the magnetic transitions, which are a consequence of the flat-ish bands of Co. Moreover, we have shown that the shorter As-As in CaCo2As2, as compaied to the P-P distance in LaCo2P2, enhances electron hopping between the pnictogen and transition metal sites and, hence, favors the AFM ordering in CaCo2As2. Our findings deepen the fundamental understanding of the electronic band structure of a large family of compounds that host a variety of collective ground states and provide evidence that the structural, electronic, and magnetic properties are strongly entangled.
Acknowledgments